# Enhancing Multimodal Large Language Models with Vision Detection Models: An Empirical Study

### v2

@@ Abstract
Despite the impressive capabilities of Multimodal Large Language Models (MLLMs) in integrating text and image modalities, challenges remain in accurately interpreting detailed visual elements. This paper presents an empirical study on enhancing MLLMs with state-of-the-art (SOTA) object detection and Optical Character Recognition (OCR) models to improve fine-grained understanding and reduce hallucination in responses. We investigate the embedding-based infusion of textual detection information, the impact of such infusion on MLLMs’ original abilities, and the interchangeability of detection models. We conduct systematic and extensive experiments with representative models such as LLaVA-1.5, DINO, PaddleOCRv2, and Grounding DINO, revealing that our simple yet general approach not only refines MLLMs’ performance in fine-grained visual tasks but also maintains their original strengths. Notably, the enhanced LLaVA-1.5 outperforms its original 7B/13B models on all 10 benchmarks, achieving an improvement of up to 12.5% on the normalized average score. We release our codes to facilitate further exploration into the fine-grained multimodal capabilities of MLLMs.
@@ 1 Introduction
The advent of large language models (LLMs) has marked a new epoch in natural language processing [3, 47], laying the groundwork for the evolution of Multimodal Large Language Models (MLLMs) that blend linguistic and visual understanding. Pioneers such as GPT4V have led the charge, showcasing remarkable proficiency in numerous tasks [53]. Despite their successes, a gap remains in these models’ ability to discern and accurately respond to queries about fine details within images [12]. This shortfall is particularly evident when MLLMs generate responses that are coherent yet misaligned with the image content, a phenomenon often referred to as “hallucination” [26, 16].
Drawing parallels to human communication, where specificity in referencing objects is key to clarity, MLLMs stand to gain from improved precision in identifying objects and texts within images. To this end, one promising yet under-explored research direction is to leverage state-of-the-art (SOTA) object detection and Optical Character Recognition (OCR) technologies to alleviate hallucination.
The prowess of contemporary object detection and OCR models in precisely delineating objects and texts within images is well-established [62, 34]. In this paper, we hypothesize that their outputs can serve as the foundation for improving MLLMs’ fine-grained understanding of images, and investigate how to effectively infuse this information. Based on the experimental analysis of representative SOTA models including LLaVA-1.5 [30], DINO [59], PaddleOCRv2 [11] and Grounding DINO [32] in the main page, as well as the experimental analysis of Qwen-VL [2] and YOLOv8 [20] in the Appendix, we systematically address the following issues:
(1) What if we directly input detection information into MLLMs? A straightforward method to enhance MLLMs with detection information is to feed the textual representation of detected targets directly into the model without additional training, leveraging the pre-existing language understanding abilities of the original MLLM. Our empirical study first focuses on this approach. Interestingly, the results reveal that while the 7B variant of LLaVA-1.5 shows modest gains, the 13B variant deteriorates, underscoring the need for additional model training.
(2) Does training with detection information harm MLLM’s original ability? As applying in-context learning on MLLMs results in excessively long input sequences, we opt to employ a training-based methodology to instruct the model in the recognition of detection information instead. We empirically show that retraining LLaVA-1.5 with detection information enhances its ability to understand detection results, yet dependency on this new information can overshadow MLLM’s original ability to leverage ViT [10] features, resulting in a decline in MLLM’s abilities outside of fine-grained understanding. Then, we further explore suitable fine-tuning strategies and show how to effectively harmonize these impacted abilities.
(3) What if we replace the detection models? The characteristics and performance of leveraged detection models impact the effectiveness of the enhanced MLLM. We examine the replacement of DINO with an open-set object detection model called GroundingDINO [32]. The result shows further enhancement on the MLLM, with new capabilities to dynamically align object detection with the context of user queries during inference. The result also underscores the robustness of the studied infusion approach, as the enhanced MLLM remains functional and effective following the replacement of detection models, leaving great potential for future stronger detection models.
As a result, our experiments affirm the value of detection model infusion and provide several practical insights, with our enhanced LLaVA-1.5 significantly outperforming its original models across all 10 benchmarks, making a notable improvement in multimodal understanding. As shown in Figure 1, the enhanced model effectively reduces hallucination and showcases enhanced efficacy in counting and localization tasks. For OCR tasks, it can also produce more accurate responses.
@@
Figure 1: Examples on which LLaVA-1.5-13B fails while our model with detection information (FTBI-13B) succeeds. “Detection” indicates processed detection information from OD/OCR models. More examples are provided in Figure 5 of Appendix A.1.
Our contributions can be summarized as follows:
  • We conduct a thorough empirical investigation for the integration of both object detection and OCR models into an MLLM, with a particular focus on infusion strategies. Furthermore, we explore the replacement of an open-set detector to enable question-driven detection and validate the sustained efficacy of our method following the substitution of detection models.
  • Built upon the insights gained, we develop a simple yet general approach for fine-grained image understanding, which elevates LLaVA-1.5 with 12.5% and 11.5% overall improvements for 7B and 13B size respectively, across 10 comprehensive MLLM benchmarks.
  • Our code is publicly available at anonymous link to facilitate further research, which we hope to help pave the way for systems that engage in more nuanced and accurate multimodal dialogue.
2 Related Works
@@ 2.1 Multimodal Large Language Models (MLLMs)
Large Language Models (LLMs) are primarily tailored for text-based tasks [61]. To integrate visual information, modality bridging modules have been developed to reconcile the representational disparities between text and images [55]. Commonly, there are three kinds of them:
Firstly, learnable queries are used to distill information from image features. For example, Flamingo’s [1] perceiver resampler and IDEFICS’s [19, 24] similar modules employ learnable queries to extract ViT features. BLIP-2 [27] uses learnable queries in conjunction with a Q-Former module, while Qwen-VL [2] compresses visual features into a sequence of fixed length using a cross-attention layer. Secondly, projection-based interfaces are employed to bridge modalities. Noteworthy models like LLaVA [31] and LLaVA-1.5 [30] use simple linear layers for mapping image features into the text semantic space. Lastly, parameter-efficient tuning modules are used. LLaMA-Adapter [60, 13] introduces self-attention layers with zero gating, and LaVIN [37] employs modality-specific adapters.
In the main page, we present experimental results for LLaVA-1.5 because its architecture aligns with the majority of current SOTA MLLMs, making it an appropriate representation. More detailed discussion and empirical support are provided in Appendix D.1, where we also conduct experiments on another MLLM, Qwen-VL, and obtain results with similar trends to verify our conclusions.
@@ 2.2 Enhancing Detection Capabilities for MLLMs
For models designed for image tasks, the ability to detect and localize targets is crucial. To improve MLLMs’ detection capabilities, typically three categories of methods are explored:
@@
The first category is to expand the dataset with existing object detection or OCR data. InstructBLIP [8] uses data from 26 datasets spanning 11 tasks, including OCR data. Shikra [6] integrates object detection dataset RefCOCO [21] and PointQA [38], and generates more related data with GPT-4 and Flickr30K [42]. ASM [48] introduces 1B region-text pairs. LLaVA and SPHINX [29] compile hybrid instruction fine-tuning datasets including object detection dataset RefCOCO, VG [22], and OCR dataset OCRVQA [40]. PINK [52] employs a bootstrapping method to cover a range of referential comprehension datasets.
The second category is to restructure the image encoder to extract fine-grained features. LION [5] introduces a module called Vision Aggregator for feature aggregation. Honeybee [4] uses a deformable attention-based abstractor with initialized reference points for fine-grained detail capture. UReader [54] employs a shape-adaptive cropping module and learnable queries to process local image features. Vary [51] develops a dedicated image encoder for text recognition. Although these approaches can enhance the detection capabilities of MLLMs, they usually require a considerable volume of data to train the supplementary modules.
@@
The third category is to introduce detection models into MLLMs’ output end to train MLLMs or perform detection tasks. UNIFIED-IO [36, 35] unifies image, text, and detection features into discrete tokens and trains an end-to-end MLLM capable of detecting. ContextDET [57] trains a visual decoder for predicting bounding boxes with contextual LLM tokens. Lenna [50], Lisa [23], and Next-chat [58] introduce an additional token to prompt the detector to identify targets. Similar to the second category, these approaches also require a substantial amount of data to achieve feature alignment.
Distinct from them, we opt not to expand datasets or introduce new architectures. Instead, we use the same datasets and architectures of the target MLLM to be infused, and employ off-the-shelf detection models to infuse textual detection information into the input end of the MLLM. Doing so allows MLLMs to achieve strong performance with minimal training effort as shown in our experiments.
3 Investigation Methodology for Detection Models Infusion
3.1 Motivational Observation
Existing MLLMs often encounter difficulties in accurately detecting targets. Consider LLaVA-1.5 as an example; despite its exceptional performance across various vision-language tasks, it’s not infallible. As shown in Figure 1, LLaVA-1.5 miscounts a herd of sheep, which points to a limitation in its object-counting capability. Besides, LLaVA-1.5 fails to detect a pedestrian on the street who is partially obscured by a utility pole, underscoring a weakness in its object localization proficiency. In another scenario, LLaVA-1.5 incorrectly recognizes the license plate number “87025” as “547”, revealing a shortcoming in its text recognition functionality.
By contrast, SOTA object detection and OCR models demonstrate superior performance in these tasks. As depicted in Figure 1, detection models yield precise “Detection” annotations. Motivated by the limitations of current MLLMs, we aim to explore the infusion of off-the-shelf detection models into MLLMs. The following section details our empirical investigation into effective fusion strategies.
@@ 3.2 Studied Model Architecture
Text-Based Detection Information Construction.
[…]
With OCR models, we can extract textual content within images along with their corresponding positional information. In the main page, we adopt PaddleOCRv2 [11] as a representative, a lightweight SOTA OCR system. Similar to object detection, we extract the textual content and corresponding central coordinates from OCR results, process it into text form, and then prepend an instruction sentence to obtain the final sentence, e.g., “Here are the central coordinates of certain texts in this image: ‘Birthday’[0.41, 0.85], ‘YEARS’[0.11, 0.34].
Specific examples with more images are provided in Appendix A.2. Besides, in Appendix B.1, we conduct statistical analyses on the length of extracted texts, showing that this simple-to-implement constructing method effectively expresses useful information as well as compress the length.
The MLLM Architecture.
Figure 2 illustrates the overall architecture of our studied models in different settings, taking the LLaVA-1.5 as an example of the target MMLM to be infused. Firstly, the CLIP-ViT-L-336px [43] is used to extract image-level features and a two-layer MLP is employed to map these features into the text semantic space. Subsequently, we separately use DINO and PaddleOCRv2 to obtain object detection and OCR results. These results are then transformed into sentences using the methods mentioned above and converted into text features using the embedding layers of the backbone LLM. Next, we concatenate the image-level features and the detection features before feeding them into the backbone LLM. As a result, the MLLM can simultaneously obtain both the overall image information and the fine-grained image details during training and inference.
Figure 2: The illustration of investigated MLLM architectures with different strategies to infuse detection information. “(LLaVA-1.5)” denotes module initialization with weights from LLaVA-1.5.
@@ 3.3 Studied Infusion Strategies
Besides the model architecture, the training strategy plays a vital role in the detection model infusion. We design a series of experiments to identify effective fusion strategies in terms of model training as follows. We provide a concise introduction here and more implementation details in Appendix B.
Training-free Infusion (TFI) For the first fusion strategy, we directly feed the detection information into the target MLLM (LLaVA-1.5 here) without any additional training. As shown in Figure 2(a), we use the same model structure and parameter as the target MLLM, with the only distinction being the supplementary input of our detection information (described in Section 3.2).
Retraining Based Infusion (RBI) For the second fusion strategy, we train our model from scratch using the target MLLM’s training pipeline. As shown in Figure 2(b), we first initialize the MLP module and pretrain it with the LCS-558K datasets [30]. Subsequently, we introduce LoRA [15] modules into the backbone LLM, Vicuna-1.5 [7]. After that, we train the LoRA modules and the MLP module during the instruction fine-tuning process with the target MLLM’s original instruction-following dataset, whose details are provided in Appendix B.2. Throughout the entire training process, we continuously infuse the detection information.
Fine-tuning Based Infusion (FTBI) For the third fusion strategy, we conduct fine-tuning on the target MLLM. As shown in Figure 2(c), we freeze the weights of both the MLP module and backbone LLM of the target MLLM. Following this, we introduce LoRA modules to the LLM and train the LoRA module for a single epoch with the target MLLM’s original instruction-following dataset, concurrently infusing the detection information.

### v3

# From Training-Free to Adaptive: Empirical Insights into MLLMs’ Understanding of Detection Information

@@ Abstract
Despite the impressive capabilities of Multimodal Large Language Models (MLLMs) in integrating text and image modalities, challenges remain in accurately interpreting detailed visual elements. Fortunately, vision detection models have shown superior performance in recognizing fine-grained image details, leading to their increased deployment by researchers to enhance the ability of MLLMs. Among the feasible strategies, infusing detection information in text format is easy to use and effective. However, most studies apply this method in a training-free manner. There is limited research on the effects of adaptive training, which has great potential for helping LLMs better comprehend the special input and discard irrelevant information. In this paper, we address the key research question: How does training influence MLLMs’ understanding of infused textual detection information? We systematically conduct experiments with numerous representative models to explore the performance implications of training-free, retraining, and fine-tuning strategies when infusing textual detection information into MLLMs. Additionally, we investigate the impact of training on the original abilities of MLLMs, as well as the interchangeability of detection models. We find that fine-tuning the pre-trained MLLM to adapt to textual detection information yields better results compared to the training-free strategy and the retraining strategy, with the fine-tuned MLLM outperforms the training-free MLLM by 6.71% across 10 widely recognized benchmarks. Besides, we find that fine-tuning allows the MLLM to maintain performance improvements even after replacing the deployed detection models, which means that it enables the MLLM to better understand the specially formatted textual information. We release our codes to facilitate further exploration into the fusion strategies of vision detection models and improving the fine-grained multimodal capabilities of MLLMs.
@@ 1 Introduction
The advent of large language models (LLMs) has marked a transformative era in natural language processing (Brown et al., 2020; Touvron et al., 2023), paving the way for the development of Multimodal Large Language Models (MLLMs) that blend linguistic and visual understanding. Pioneers such as GPT-4V have demonstrated remarkable proficiency across numerous tasks (Yang et al., 2023). However, a notable gap remains in these models’ ability to accurately discern and recognize fine details within images (Fu et al., 2023). This limitation is particularly evident when MLLMs generate coherent yet misaligned responses with the image content, a phenomenon often referred to as “hallucination” (Li et al., 2023b; Huang et al., 2023).
Current advancements in object detection and optical character recognition (OCR) models have established their effectiveness in identifying objects and text within images (Zou et al., 2023; Liu et al., 2024c). Consequently, researchers have increasingly deployed vision detection models to assist MLLMs in recognizing fine-grained visual elements. A popular approach involves converting the outputs of vision detection models into textual descriptions, which are then supplied to the backbone LLM, thereby enhancing the MLLM’s performance in visual tasks. This fusion strategy is both straightforward and effective.
Nonetheless, the majority of existing research has primarily focused on training-free methods to directly apply the textual detection information  [^1] To maintain brevity, we refer to “textual detection information” as the information output by vision detection models in textual format.. Little exploration has been conducted into adaptive training methods, which have great potential to enhance LLMs’ comprehension of specially formatted textual content, enabling them to intentionally discard irrelevant information and generate more pertinent responses (Zhang et al., 2024b; Cabessa et al., 2024). This highlights the need for a systematic investigation, particularly concerning the core research question: Can adaptive training further enhance MLLMs’ performance beyond what is achievable through training-free integration of textual detection information?
@@
Figure 1: Examples where LLaVA-1.5-13B fails, while the model infused with textual detection information (FTBI-13B) succeeds. “Detection” refers to processed detection information from OD/OCR models. Additional examples are provided in Figure 5 of Appendix A.1.
To provide insights into how training impacts the infusion of textual detection information into MLLMs, we investigate training-free, retraining, and fine-tuning strategies for this fusion method. Additionally, we examine how training influences the original image understanding capabilities of MLLMs and the interchangeability of deployed detection models. Based on the experimental analysis encompassing representative advanced models, including LLaVA-1.5 (Liu et al., 2023a), DINO (Zhang et al., 2022), PaddleOCRv2 (Du et al., 2021), and Grounding DINO (Liu et al., 2023c), alongside Qwen-VL (Bai et al., 2023) and YOLOv8 (Jocher et al., 2023) in the appendix, we systematically uncover the following key insights:
(1) The fine-tuning strategy yields better results than both the training-free and retraining strategies.
[…]
(2) Retraining with textual detection information impairs MLLMs’ original image comprehension abilities.
Most advanced MLLMs employ an image encoder to generate image features, and their ability to understand these features is crucial for effective multimodal understanding. Our experiments reveal that retraining the MLLM with textual detection information detrimentally affects its ability to interpret the features from its image encoder. In contrast, the fine-tuning strategy does not run into this problem.
(3) Fine-tuning allows the MLLM to retain performance improvements upon replacing the deployed detection model.
The characteristics and performance of the deployed detection models significantly influence the enhanced MLLM’s effectiveness. Based on the fine-tuning strategy, we examine replacing a closed-set detector with an open-set detector. The results demonstrate further enhancement in MLLM performance, enabling dynamic object detection following the context of user queries during inference. Additionally, we find that the fine-tuned MLLM maintains its training benefits and can still effectively discard irrelevant information even after the model replacement.
To summarize, our work contributes comprehensive empirical evidence and practical insights into the effects of various training strategies for infusing textual detection information into MLLMs. It identifies a significant gap between the use of adaptive training and training-free methods, highlighting the potential of adaptive strategies and demonstrating their feasibility through systematic investigation. Our code is publicly available at anonymous link to facilitate further research and pave the way for systems that engage in more nuanced and accurate multimodal dialogue.
2 Background and Motivation
@@ 2.1 Multimodal Large Language Models (MLLMs)
Linking Text and Vision Information. Large Language Models (LLMs) are primarily designed for text-based tasks (Zhao et al., 2023). To incorporate image processing capabilities, modality bridging modules have been developed to reconcile the representation differences between text and images (Yin et al., 2023; Qin et al., 2024). Generally, these methods can be categorized into three types:
(1) Learnable queries are used to distill information from image features. For instance, Flamingo (Alayrac et al., 2022) employs a perceiver resampler, and IDEFICS (Hugo et al., 2023; Laurençon et al., 2024) uses similar modules to extract features from Vision Transformers (ViT) (Dosovitskiy et al., 2020). BLIP-2 (Li et al., 2023c) utilizes learnable queries alongside a Q-Former module, while Qwen-VL (Bai et al., 2023) compresses visual features into sequences of fixed length using cross-attention layers. (2) Projection-based interfaces bridge modalities with straightforward techniques. Notable examples include LLaVA (Liu et al., 2023b; a; 2024a) and MGM (Li et al., 2023d), which utilize simple linear layers to map image features into the text semantic space. (3) Parameter-efficient tuning modules are utilized to fine-tune MLLMs for image feature comprehension. For example, LLaMA-Adapter (Zhang et al., 2023; Gao et al., 2023) introduces self-attention layers with zero gating for fine-tuning, and LaVIN (Luo et al., 2023) employs modality-specific adapters.
Why Incorporating Detection Models into MLLMs? Existing MLLMs often struggle to accurately detect fine-grained targets. For example, in Figure 1, LLaVA-1.5 miscounts a herd of sheep, indicating a limitation in its object-counting capability. Additionally, it fails to detect a pedestrian who is partially obscured by a utility pole, highlighting a weakness in its object localization ability. In another scenario, LLaVA-1.5 incorrectly recognizes the license plate number “87025” as “547”, revealing a shortcoming in its text recognition ability. By contrast, SOTA object detection and OCR models demonstrate superior performance on detection and recognition tasks, which has led many researchers to explore the application of detection models within the realm of MLLM research.
@@ 2.2 Enhancing Detection Capabilities for MLLMs
Existing Methods for Detection Capabilities Enhancement. Various strategies have been explored to enable MLLMs aware of image details, generally classified into four types:
@@
(1) Expanding datasets with existing object detection or OCR data: InstructBLIP (Dai et al., 2023) utilizes data from 26 datasets across 11 tasks, including OCR data. ASM (Wang et al., 2023a) introduces 1 billion region-text pairs. LLaVA and SPHINX (Lin et al., 2023) compile hybrid instruction fine-tuning datasets, incorporating object detection datasets like VG (Krishna et al., 2017) and the OCR dataset OCRVQA (Mishra et al., 2019). PINK (Xuan et al., 2023) employs a bootstrapping method to cover diverse referential comprehension datasets. MiniGPT4-v2 (Chen et al., 2023b), VisionLLM (Wang et al., 2024b), and Shikra (Chen et al., 2023c) integrate object detection datasets, such as RefCOCO (Kazemzadeh et al., 2014), PointQA (Mani et al., 2020), and Flickr30K (Plummer et al., 2015), while introducing special detection tokens like “det” to guide downstream tasks (further details in Appendix D.7).
(2) Restructuring the image encoder to extract fine-grained features: LION (Chen et al., 2023a) introduces a Vision Aggregator module for feature aggregation, while Honeybee (Cha et al., 2023) employs a deformable attention-based abstractor for capturing fine details. UReader (Ye et al., 2023) utilizes a shape-adaptive cropping module to process local image features, and Vary (Wei et al., 2023b) develops a dedicated image encoder for text recognition. Eagle (Shi et al., 2024) aligns features from various visual experts, concatenating them as input for the MLLM. Mova (Zong et al., 2024) introduces the MoV-Adapter, which extracts and fuses task-specific knowledge.
@@
(3) Integrating pre-trained detection models into MLLMs’ output end to train MLLMs or perform detection tasks: UNIFIED-IO (Lu et al., 2022; 2023) unifies image, text, and detection features into discrete tokens and trains an end-to-end MLLM capable of detecting. ContextDET (Zang et al., 2023) trains a visual decoder for bounding box prediction using contextual LLM tokens. Lenna (Wei et al., 2023a), Lisa (Lai et al., 2023), and Next-chat (Zhang et al., 2024a) introduce additional tokens to prompt detectors for target identification.
(4) Converting detection model outputs into text and using it as supplementary input for LLMs: GLEE (Wu et al., 2024) builds on LISA (Lai et al., 2023) to generate SEG tokens for targeted segmentation, enhancing performance by feeding textual object queries into the backbone LLM. P2G (Chen et al., 2024d), Moai (Lee et al., 2024), and IVE (He et al., 2024) employ detection agents to generate textual grounding clues for improved reasoning. Power-LLaVA (Wang et al., 2024a) utilizes an object detector to produce textual class and location information to assist the MLLM in generating high-quality outputs. VLPrompt (Zhou et al., 2023) leverages an object detector to generate target names and infer relationships, thereby aiding MLLMs in reasoning tasks.
Why Adaptive Training with Textual Detection Information? Although methods in the second and third categories can improve the detection capabilities of MLLMs, they typically require substantial datasets to train the restructured image encoders or achieve feature alignment. In contrast, text-based methods are simpler, necessitate less extensive training data for the newly built detection modules, and still deliver commendable results. Thus, the fourth type of method is likely to be more frequently employed in practical applications.
While most research in this category has concentrated on training-free strategies for infusing textual detection information, we note relevant developments in pure-text LLMs. Zhang et al. (2024b) propose leveraging Retrieval-Augmented Generation (RAG, Gao et al. (2024)) during fine-tuning to help LLMs discard redundant information from augmented text. Additionally, Cabessa et al. (2024) suggest that infusing well-crafted textual features during fine-tuning can enhance LLMs’ comprehension of specially formatted inputs. They primarily use a small amount of data to adaptively train LLMs for comprehending specially formatted text, yielding excellent results.
This leads us to an important question: Since the infusion of textual detection information already performs well without training, could this fusion method achieve even better outcomes with appropriate training? Our work aims to address this question by utilizing the original training data of the studied MLLMs, which is limited in quantity but high in quality, to conduct adaptive training for the infusion of textual detection information into MLLMs.
3 Investigation Methodology for the Infusion of Textual Detection Information
3.1 Text-Based Detection Information Construction
Following the typical practice of many previous studies (Wang et al., 2024a; Chen et al., 2024d; Zhou et al., 2023; Liu et al., 2024b), we first need to convert the output of object detection models and OCR models into specially formatted text.
Object Detection Information.
[…]
OCR Information.
With OCR models, we can extract textual content within images along with their positional information. In the main page, we adopt PaddleOCRv2 (Du et al., 2021) as a representative, a lightweight SOTA OCR system. Similar to what we’ve done for object detection information, we extract the textual content and corresponding central coordinates from OCR results, process them into text form, and then prepend an instruction sentence to obtain the final sentence, e.g., “Here are the central coordinates of certain texts in this image: ‘Birthday’[0.41, 0.85], ‘YEARS’[0.11, 0.34].
Examples.
Specific examples with images are provided in Appendix A.2. In Appendix B.1, we conduct statistical analyses on the length of processed texts, showing that this simple-to-implement constructing method effectively expresses useful information as well as compress the length.
@@ 3.2 Studied Model Architecture
Specifically, Figure 2 illustrates the overall architecture of the studied MLLM in different training strategies, taking LLaVA-1.5 as an example.  [^2] On the main page, we mainly focus on LLaVA-1.5 due to its architectural alignment with many leading MLLMs, making it a representative choice. For more detailed discussions and empirical evidence, please refer to Appendix D.1, where we also present findings from experiments on another MLLM, Qwen-VL, which yield similar trends to corroborate our conclusions. Firstly, the CLIP-ViT-L-336px (Radford et al., 2021) is used to extract image-level features and a two-layer MLP is employed to align these features with text. Subsequently, we separately use DINO and PaddleOCRv2 for object detection and OCR. The results are then converted into sentences using the aforementioned methods and transformed into text features using the embedding layers of the backbone LLM. Next, we concatenate the image-level features and the detection features and input them into the backbone LLM. As a result, the MLLM can simultaneously obtain both the overall image information and the fine-grained image details during training and inference.
Figure 2: The studied MLLM architectures with different training strategies for infusing textual detection information. “(LLaVA-1.5)” denotes module initialization with weights from LLaVA-1.5.
@@ 3.3 Studied Infusion Strategies
We systematically design three training strategies for the infusion of textual detection information, using LLaVA-1.5 as a representative. We provide more implementation details in Appendix B.
Training-free Infusion (TFI). For the first strategy, we directly feed the textual detection information into the MLLM without any additional training. As shown in Figure 2(a), we use the same model structure and parameter as the studied MLLM, with the only distinction being the supplementary input of the textual detection information.
Retraining Based Infusion (RBI). For the second strategy, we train the model from scratch using the studied MLLM’s training pipeline. As shown in Figure 2(b), we first initialize the MLP module and pre-train it with the studied MLLM’s original pre-training dataset. Subsequently, we introduce LoRA (Hu et al., 2021) modules into the backbone LLM, Vicuna-1.5 (Chiang et al., 2023). After that, we train the LoRA modules and the MLP module during the instruction tuning process with the studied MLLM’s original instruction-following dataset, whose details are provided in Appendix B.2. Throughout the entire training process, we continuously infuse the textual detection information.
Fine-tuning Based Infusion (FTBI). For the third strategy, we conduct fine-tuning on a well-trained MLLM. As shown in Figure 2(c), we freeze the weights of both the MLP module and the backbone LLM of the pre-trained MLLM. Following this, we introduce LoRA modules to the LLM and train the LoRA modules for a single epoch with the studied MLLM’s original instruction-following dataset, concurrently infusing the textual detection information.